a chemist needs 9 liters of a 10% alcohol solution but has only a 15% alcohol sulotion. how many liters each of the 15% solution and water should he mix to make the desired 9 liters of 10% solition
Answer by lwsshak3(11628)

a chemist needs 9 liters of a 10% alcohol solution but has only a 15% alcohol sulotion. how many liters each of the 15% solution and water should he mix to make the desired 9 liters of 10% solition
***
let x=amt of water to mix
9-x=amt of 15% solution to mix
..
15%(9-x)=10%*9
1.35-.15x=10%*9
.15x=1.35-.90
.15x=.45
x=3
9-x=6
amt of water to mix=3 liters
amt of 15% solution to mix=6 liters